# Optimal structure learning and conditional independence testing

### v1

@@ 1.1 Contributions
    We establish a connection (Theorem 3.1) between the statistical complexity of conditional independence testing and structure learning problems. We show that the optimal sample complexity of learning any poly-forest is
@@
    when the optimal testing radius of the corresponding CI testing problem is $n^{-1/\alpha}$ for some $\alpha>0$ depending on the modeling setup, and $c$ is a parameter that captures the signal strength (cf. Section 2).
@@ 1.2 Related work
CI testing forms a crucial building block in machine learning and reasoning tasks. Given a triplet of random variables $(X,Y,Z)$, the problem asks whether $X$ is conditionally independent of $Y$ given $Z$. Numerous methods (e.g. Dawid, 1979; Fukumizu et al., 2007; Ramsey, 2014; Li and Fan, 2020) have been developed to address this fundamental problem. One prominent class of methods relies on measuring the distance between the conditional distribution $P(X,Y\,|\,Z)$ and the product of the marginals $P(X\,|\,Z)P(Y\,|\,Z)$. For Gaussian variables, the problem reduces to testing whether the partial correlation is zero (Fisher, 1915). For discrete variables, hypothesis tests based on chi-squared statistics are widely employed Ireland and Kullback (1968); Darroch et al. (1980). Recent advancements have also explored kernel-based methods Gretton et al. (2007); Zhang et al. (2011) and information-theoretic measures such as conditional mutual information (Runge, 2018; Berrett and Samworth, 2019) that can capture complex nonlinear dependencies in a nonparametric fashion.
@@
Learning the structure of graphical models from observational data is a fundamental problem in machine learning. For undirected graphical models, the problem reduces to support recovery of the precision matrix (MEINSHAUSEN and BÜHLMANN, 2006; Friedman et al., 2008; Cai et al., 2011; Liu et al., 2009). For DAG learning, approaches can be broadly categorized into three main paradigms: score-based methods (e.g. greedy equivalence search) (Chickering, 2002; Nandy et al., 2018), constraint-based methods (e.g. PC algorithm) (Spirtes and Glymour, 1991; Friedman et al., 2013), and methods that leverage specific distributional assumptions (Shimizu et al., 2006; Hoyer et al., 2008; Peters et al., 2011, 2014; Peters and Bühlmann, 2014). Especially, constraint-based methods rely on valid CI tests and the faithfulness assumption to operate (Kalisch and Bühlman, 2007; Marx et al., 2021), thus development of reliable and efficient CI tests directly impacts the performance of these structure learning methods. For tree-structured models, the Chow-Liu algorithm (Chow and Liu, 1968; Chow and Wagner, 1973) provides an efficient method to find the optimal undirected tree structure by constructing a maximum weight spanning tree based on pairwise mutual information. Learning poly-trees/forests is generally more complex while more tractable than learning general DAGs REBANE (1987); Srebro (2003); Tan et al. (2010, 2011). Recent developments include (Gao and Aragam, 2021; Azadkia et al., 2021; Jakobsen et al., 2022). Furthermore, learning poly-trees/forests is relevant in various applications, including reasoning and density estimation (Kim and Pearl, 1983; Liu et al., 2011).
@@ 2.1 Measuring dependence in general models
Let $(X,Y,Z)\sim p$ be a triplet of random variables subject to some distribution $p$. A (conditional) dependence measure $m(X;Y\,|\,Z)$ is a functional of $p$ such that (1) $m(X,Y\,|\,Z)\geq 0$; and (2) $m(X;Y\,|\,Z)=0$ if and only if $X\perp\!\!\!\perp Y\,|\,Z$ under $p$.
If $Z=\emptyset$, we simplify the notation by writing the marginal dependence as $m(X;Y)=m(X;Y\,|\,\emptyset)$. This general measure of dependence is used to simplify our main theorem statement; in our examples (Sections 4-5), we specify the dependence measure as the usual correlation coefficient for Gaussian distributions, total variation with the product of marginals for Bernoulli distributions, and total variation distance to the nearest independent instance for nonparametric continuous distributions. Using this dependence measure, we can now define strong tree-faithfulness in a generic form.
@@
    For any two nodes connected $j-k$, we have $m(X_{k};X_{j}\,|\,X_{\ell})\geq c$ for $\ell\in V\cup\{\emptyset\}\setminus\{k,j\}$;
@@ 2.2 CI testing and structure learning
Let $\mathcal{P}$ be all continuous distributions supported on $[0,1]^{3}$ that admit Lipschitz continuous densities $p$. For each distribution $p\in\mathcal{P}$ over a triplet of variables $(X,Y,Z)$, we measure the dependence between $X$ and $Y$ given $Z$ by the total variation distance $\inf_{q\in\mathcal{Q}}\|p-q\|_{1}$, where $\mathcal{Q}$ is the set of all conditionally independent distributions in $\mathcal{P}$. This serves as a valid dependence measure $m$ for nonparametric distributions. The nonparametric CI testing problem asks for a test to distinguish the following two hypotheses:
@@
where $p\in\mathcal{P}$. A conditional independence test $\psi$ is a function of data $\{X^{(i)},Y^{(i)},Z^{(i)}\}_{i=1}^{n}\mapsto\{0,1\}$; i.e. $\psi=0$ indicates $X\perp\!\!\!\perp Y\,|\,Z$. The minimax optimal testing radius of $\mathcal{C}(\mathcal{P},m,c)$ is the infimum of $c=c_{n}>0$ in terms of sample size $n$ such that the following holds:
@@
The number $1/10$ here can be replaced with any small constant less than one. For a certain CI testing problem, one needs to specify the model class $\mathcal{P}$, and the dependence measure $m$. Consequently, the hypothesis classes $\mathcal{H}_{0},\mathcal{H}_{1}$ are determined. In Example 1, $\mathcal{P}$ includes all the Lipschitz densities over $[0,1]^{3}$ and $m$ is given by the total variational distance. One goal of studying $\mathcal{C}(\mathcal{P},m,c)$ is to derive the minimax testing radius $c$ such that the conditionally dependent and independent instances can be distinguished accurately. Crucially, the hypotheses $\mathcal{H}_{0}\text{ and }\mathcal{H}_{1}$, together with the dependence measure $m$, can be used to characterize the $c$-strong tree-faithfulness. We continue to introduce the poly-forest learning problem using this set-up.
@@
The poly-forest learning problem $\mathcal{F}(\mathcal{P},m,c)$ is to learn the Markov equivalence class of a poly-forest $G\in\mathcal{T}$ given i.i.d. samples from a distribution $p$ that is Markov and $c$-strong tree-faithful to $G$ with respect to the dependence measure $m$, and for any triplet of nodes $(j,k,\ell)$, the marginal $p(X_{j},X_{k},X_{\ell})$ comes from either $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$. The optimal sample complexity of $\mathcal{F}(\mathcal{P},m,c)$ refers to the sufficient and necessary conditions on sample size $n$ in terms of the number of nodes $d$ and the strong tree-faithfulness parameter $c$, such that for some estimator $\widehat{G}$ and some constant $\delta\in(0,1)$, the following holds:
@@ 5 Application to nonparametric models
We adopt the framework in Neykov et al. (2021) and provide the essential elements in Appendix B.3, see details therein. As alluded to in Example 1, we consider all continuous distributions over $[0,1]^{3}$, which admit continuous densities $p(X,Y,Z)$. We measure the dependence using
@@
[…]
@@ 6 Experiments
In Figure 1, we present the structure Hamming distance (SHD) between the estimated graph and the ground truth against sample size $n$ for various settings of number of nodes $d=\{20,40,60,80,100\}$. Each setting is averaged over $N=20$ replications. The results show as the sample size increases, SHD consistently decreases for graphs with 20 to 100 nodes. A lower SHD means the output graph is closer to the truth, thus more accurate structure learning. The effect of dimensionality is illustrated by lines with different colors. Overall, our simulations demonstrate that combining a powerful conditional independence test $\psi$ with PC-tree algorithm allows for consistent and accurate poly-forest learning. Full experiment details on implementations and how the data are simulated, and additional experiment results are postponed to Appendix H.
@@ 7 Conclusion
In this paper, we study the minimax optimality of structure learning and conditional independence testing. We establish a connection between CI testing and structure learning by showing the optimality conditions for CI testing translate directly into the optimality of poly-forest learning, which can be achieved by an efficient constraint-based algorithm with the optimal CI test as input. The generic theoretical results are demonstrated using three representative distribution families. This finding highlights the central role of CI testing in structure learning in the view of statistical sample efficiency.
One interesting direction for future research is to extend the results beyond poly-forests to general DAGs. In such settings, constraint-based methods — such as the PC algorithm — are applicable and rely solely on CI testing to operate. We anticipate that a similar statistical connection between CI testing and structure learning can be established for general DAGs, though additional structural parameters, such as the maximum in-degree, are likely to play a key role in characterizing the corresponding optimal sample complexity.
@@
      If all tests reject, then $\widehat{E}\leftarrow\widehat{E}\cup\{j-k\}$.
@@
    • R2 Orient $j-k$ into $j\rightarrow k$ whenever there is a chain $j\rightarrow\ell\rightarrow k$
    • R3 Orient $j-k$ into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow k$ and $j-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
    • R4 Orient $j-k$ into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow i$ and $\ell-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
@@ B.3 Nonparametric continuous distribution (Section 5)
We follow the CI testing framework in Neykov et al. (2021). Consider all continuous distributions over $[0,1]^{3}$ that admit continuous densities $p(X,Y,Z)$. In addition, we impose following two smoothness condition on the densities.
@@
$\mathcal{H}_{0}^{NP}$ =$=\{p\in\mathcal{P}^{NP}:X\perp\!\!\!\perp Y\,|\,Z\text{ and }\text{satisfies Lipschitzness}\}$
@@
Correspondingly, the nonparametric poly-forest learning problem is defined, which includes all continuous distributions supported on $[0,1]^{d}$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$, denoted as $\mathcal{F}^{NP}:=\mathcal{F}(\mathcal{P}^{NP},m^{NP},c)$.
@@
Given a conditional independence testing problem $\mathcal{C}(\mathcal{P},m,c)$, if for some constants $A$ and $A^{\prime}$,
@@
$\mathbb{P}(\psi^{\prime}\text{ is incorrect})\leq\exp\Big{(}-C_{0}n/N\Big{)}\,.$
@@ D.1 Proof of upper bound for Bernoulli distribution
$|\widehat{m}^{B}-m^{B}|\leq c/2\implies\psi^{B}\text{ is correct.}$.
@@ D.2 Proof of lower bound for Bernoulli distribution
$h_{ij,m^{\prime}}(x,y)$ =$=\begin{cases}\sqrt{m^{\prime}}^{2}\widetilde{h}(m^{\prime}x-i+1,m^{\prime}y-j+1)&\forall(x,y)\in[\frac{i-1}{m},\frac{i}{m}]\otimes[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
$h_{j,m}(z)$ =$=\begin{cases}\sqrt{m}h(mz-j+1)&\forall z\in[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
[…]
@@
$p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)$ =$=\bigg{\{}\gamma_{\Delta}(x,y)$
@@
[…]
@@
We then check the TV distance between $p(x,y\,|\,z)$ and $p(x,y\,|\,z^{\prime})$:
@@
We check the Hölder smoothness of $p(x,y\,|\,z)$ in $(x,y)$. Following the proof Theorem 4.2 in Neykov et al. (2021), for any $k\leq\lfloor s\rfloor$

### v2

@@ 1.1 Contributions
    We establish a connection (Theorem 3.1) between the statistical complexity of conditional independence testing and structure learning problems. We show that the minimax optimal sample complexity of learning any poly-forest is
@@
    when the minimax optimal testing radius of the corresponding CI testing problem is $n^{-1/\alpha}$ for some $\alpha>0$ depending on the modeling setup, and $d$ is the number of nodes, $c$ is a parameter that captures the signal strength (cf. Section 2).
@@ 1.2 Related work
CI testing forms a crucial building block in machine learning and reasoning tasks. Given a triplet of random variables $(X,Y,Z)$, the problem asks whether $X$ is conditionally independent of $Y$ given $Z$. Numerous methods (e.g. Dawid, 1979; Fukumizu et al., 2007; Ramsey, 2014; Li and Fan, 2020) have been developed to address this fundamental problem. One prominent class of methods relies on measuring the distance between the conditional distribution P(X,Y|Z)P(X,Y\,|\,Z) and the product of the marginals P(X|Z)P(Y|Z)P(X\,|\,Z)P(Y\,|\,Z). For Gaussian variables, the problem reduces to testing whether the partial correlation is zero (Fisher, 1915). For discrete variables, hypothesis tests based on chi-squared statistics are widely employed Ireland and Kullback (1968); Darroch et al. (1980). Recent advancements have also explored kernel-based methods Gretton et al. (2007); Zhang et al. (2011) and information-theoretic measures such as conditional mutual information (Runge, 2018; Berrett and Samworth, 2019) that can capture complex nonlinear dependencies in a nonparametric fashion.
@@
Learning the structure of graphical models from observational data is a fundamental problem in machine learning. For undirected graphical models, the problem reduces to support recovery of the precision matrix (MEINSHAUSEN and BüUhlmannHLMANN, 2006; Friedman et al., 2008; Cai et al., 2011; Liu et al., 2009). For DAG learning, approaches can be broadly categorized into three main paradigms: score-based methods (e.g. greedy equivalence search) (Chickering, 2002; Nandy et al., 2018), constraint-based methods (e.g. PC algorithm) (Spirtes and Glymour, 1991; Friedman et al., 2013), and methods that leverage specific distributional assumptions (Shimizu et al., 2006; Hoyer et al., 2008; Peters et al., 2011, 2014; Peters and Bühlmann, 2014). Especially, constraint-based methods rely on valid CI tests and the faithfulness assumption to operate (Kalisch and Bühlman, 2007; Marx et al., 2021), thus development of reliable and efficient CI tests directly impacts the performance of these structure learning methods. For tree-structured models, the Chow-Liu algorithm (Chow and Liu, 1968; Chow and Wagner, 1973) provides an efficient method to find the optimal undirected tree structure by constructing a maximum weight spanning tree based on pairwise mutual information. Learning poly-trees/forests is generally more complex than learning undirected trees models, while remaining tractable compared to learning general DAGs (Rebane, 1987; Srebro, 2003; Tan et al., 2010, 2011). Recent developments include (Gao and Aragam, 2021; Azadkia et al., 2021; Jakobsen et al., 2022). Furthermore, learning poly-trees/forests is relevant in various applications, including reasoning and density estimation (Kim and Pearl, 1983; Liu et al., 2011).
@@ 2.1 Measuring dependence in general models
Let $(X,Y,Z)\sim p$ be a triplet of random variables subject to some distribution $p$. A (conditional) dependence measure m(X;Y|Z)m(X;Y\,|\,Z) is a functional of $p$ such that (1) $m(X,Y\,|\,Z)\geq 0$; and (2) m(X;Y|Z)=0m(X;Y\,|\,Z)=0 if and only if $X\perp\!\!\!\perp Y\,|\,Z$ under $p$.
If $Z=\emptyset$, we simplify the notation by writing the marginal dependence as $m(X;Y)=m(X;Y\,|\,\emptyset)$. This general measure of dependence is used to simplify our main theorem statement; in our examples (Sections 4-5), we specify the dependence measure as the usual correlation coefficient for Gaussian distributions, total variation with the product of marginals for Bernoulli distributions, and total variation distance to the nearest independent instance for nonparametric continuous distributions (Canonne et al., 2018; Neykov et al., 2021; Wang et al., 2024). In more general settings (e.g. user-defined models), the dependence measure can be chosen based on modeling preference, as long as there exist consistent or efficient CI tests achieving provable error guarantees, it can be embedded into our framework. Using this dependence measure, we can now define strong tree-faithfulness in a generic form.
@@
    For any two nodes connected jkj-k, we have $m(X_{k};X_{j}\,|\,X_{\ell})\geq c$ for $\ell\in V\cup\{\emptyset\}\setminus\{k,j\}$;
@@
Tree-faithfulness is a relaxed version of the general faithfulness assumption, which requires the CI relationships in data distribution to reflect the existence of edges in graph (Koller and Friedman, 2009; Wang et al., 2024).
@@ 2.2 CI testing and structure learning
Let $\mathcal{P}$ be all continuous distributions supported on $0,1]^{3}$ that admit Lipschitz continuous densities $p$. For each distribution $p\in\mathcal{P}$ over a triplet of variables $(X,Y,Z)$, we measure the dependence between $X$ and $Y$ given $Z$ by the total variation distance $\inf_{q\in\mathcal{Q}}\|p-q\|_{1}$, where $\mathcal{Q}$ is the set of all conditionally independent distributions in $\mathcal{P}$. This serves as a valid dependence measure $m$ for nonparametric distributions. The nonparametric CI testing problem asks for a test to distinguish the following two hypotheses:
@@
where $p\in\mathcal{P}$. A conditional independence test $\psi$ is a function of data $\{X^{(i)},Y^{(i)},Z^{(i)}\}_{i=1}^{n}\mapsto\{0,1\}$; i.e. $\psi=0$ indicates $X\perp\!\!\!\perp Y\,|\,Z$. The minimax optimal testing radius of $\mathcal{C}(\mathcal{P},m,c)$ is the infimum of $c=c_{n}>0$ in terms of sample size $n$ such that there exists a test whose Type-I and Type-II errors are controlled:
@@
[…]
@@
The poly-forest learning problem $\mathcal{F}(\mathcal{P},m,c)$ is to learn the Markov equivalence class of a poly-forest $G\in\mathcal{T}$ given i.i.d. samples from a distribution $p$ that is Markov and $c$-strong tree-faithful to $G$ with respect to the dependence measure $m$, and for any triplet of nodes $(j,k,\ell)$, the marginal $p(X_{j},X_{k},X_{\ell})$ comes from either $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$. The optimal sample complexity of $\mathcal{F}(\mathcal{P},m,c)$ refers to the smallest integer $n$ as sample size in terms of the number of nodes $d$ and the strong tree-faithfulness parameter $c$, such that for some estimator $\widehat{G}$ and some constant $\delta\in(0,1)$, the following holds:
@@ 5 Application to nonparametric models
We adopt the framework in Neykov et al. (2021) and provide the essential elements in Appendix B.3, see details therein. As alluded to in Example 1, we consider all continuous distributions over $0,1]^{3}$, which admit continuous densities $p(X,Y,Z)$. We measure the dependence using
@@
where $\mathcal{Q}$ is the class of continuous distributions $q(X,Y,Z)$ over $0,1]^{3}$ such that $X\perp\!\!\!\perp Y\,|\,Z$, and $\|p-q\|_{1}=\int|p(x,y,z)-q(x,y,z)|dxdydz$. This measures the distance to the closest (conditionally) independent distributions. In addition, smoothness conditions are imposed on the densities (cf. Definitions 5-6), characterized by a smoothness parameter $s$ and used to contrast with the parametric cases. Together, these defines the CI testing class $\mathcal{C}^{NP}$ and the poly-forest learning problem $\mathcal{F}^{NP}$, where $NP$ denotes “nonparametric”.
@@ 6 Experiments
In Figure 1, we present the structure Hamming distance (SHD) between the estimated graph and the ground truth against sample size $n$ for various settings of number of nodes $d=\{20,40,60,80,100\}$. Each setting is averaged over $N=50$ replications. The results show as the sample size increases, SHD consistently decreases for graphs with 20 to 100 nodes. A lower SHD means the output graph is closer to the truth, thus more accurate structure learning. The effect of dimensionality is illustrated by lines with different colors. Overall, our simulations demonstrate that combining a powerful conditional independence test $\psi$ with PC-tree algorithm allows for consistent and accurate poly-forest learning. Full experiment details on implementations and how the data are simulated, and additional experiment results are postponed to Appendix H.
@@ 7 Conclusion
In this paper, we study the minimax optimality of structure learning and conditional independence testing. We make their intuitive connection rigorous and quantitative by showing the optimality conditions for CI testing translate directly into the optimality of poly-forest learning, which can be achieved by an efficient constraint-based algorithm with the optimal CI test as input. The generic theoretical results are demonstrated using three representative distribution families. This finding highlights the central role of CI testing in structure learning in the view of statistical sample efficiency.
One interesting direction for future research is to extend the results beyond poly-forests to general DAGs. In such settings, constraint-based methods–such as the PC algorithm–are applicable and rely solely on CI testing to operate. We anticipate that a similar statistical connection between CI testing and structure learning can be established for general DAGs, though additional structural parameters, such as the maximum in-degree, are likely to play a key role in characterizing the corresponding optimal sample complexity.
@@
[…]
@@
    • R2 Orient jkj-k into $j\rightarrow k$ whenever there is a chain $j\rightarrow\ell\rightarrow k$
    • R3 Orient jkj-k into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow k$ and $j-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
    • R4 Orient jkj-k into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow i$ and $\ell-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
@@ B.3 Nonparametric continuous distribution (Section 5)
We follow the CI testing framework in Neykov et al. (2021). Consider all continuous distributions over $0,1]^{3}$ that admit continuous densities $p(X,Y,Z)$. In addition, we impose following two smoothness condition on the densities.
@@
[…]
Correspondingly, the nonparametric poly-forest learning problem is defined, which includes all continuous distributions supported on $0,1]^{d}$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$, denoted as $\mathcal{F}^{NP}:=\mathcal{F}(\mathcal{P}^{NP},m^{NP},c)$.
@@
Given a conditional independence testing problem $\mathcal{C}(\mathcal{P},m,c)$, if for some constants $A$ and $A^{\prime}$ independent with sample size $n$,
@@
$\mathbb{P}(\psi^{\prime}\text{ is incorrect})\leq\exp\Big(-C_{0}n/N\Big)\,$,
for some constant $C_{0}$ independent with $n$.
@@ D.1 Proof of upper bound for Bernoulli distribution
$|\widehat{m}^{B}-m^{B}|\leq c/2\implies\psi^{B}\text{ is correct.}$
@@ D.2 Proof of lower bound for Bernoulli distribution
$h_{ij,m^{\prime}}(x,y)$ $=\begin{cases}\sqrt{m^{\prime}}^{2}\widetilde{h}(m^{\prime}x-i+1,m^{\prime}y-j+1)&\forall(x,y)\in[\frac{i-1}{m},\frac{i}{m}]\otimes[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
$h_{j,m}(z)$ $=\begin{cases}\sqrt{m}h(mz-j+1)&\forall z\in[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
$\int\widetilde{h}(x,y)dx$ $=\sqrt{m^{\prime}}h(y)\quad\int\widetilde{h}(x,y)dy=\sqrt{m^{\prime}}h(x)\,.$
@@
$p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)$ $=\bigg\{\gamma_{\Delta}(x,y)$
@@
where $g=\int|\widetilde{h}(x,y)-\frac{1}{\sqrt{m^{\prime}}}h(x)-\frac{1}{\sqrt{m^{\prime}}}h(y)|dxdy$ being a constant. And we need for either $w=x$ or $y$,
@@
We then check the TV distance between p(x,y|z)p(x,y\,|\,z) and $p(x,y\,|\,z^{\prime})$:
@@
We check the Hölder smoothness of p(x,y|z)p(x,y\,|\,z) in $(x,y)$. Following the proof Theorem 4.2 in Neykov et al. (2021), for any $k\leq\lfloor s\rfloor$